ANALYZE AND SKETCH (SKETCHING RATIONAL FRACTIONS)

a) Y = −4/(x+1)(x^2−9)

b) Y=x^2/x^4−6x+8


1
Expert's answer
2021-03-18T05:55:20-0400

a)   Y = −4/(x+1)(x^2−9)

Y= (4x+4) (8x-36)

           4

Y= (x+1) (2x-9)

Y= (2x²-8)

 

b)   Y=x^2/x^4−6x+8

Y= x^2   

  (x^4) (-6x + 8)

Y= x

Y= (x^2) (-6x + 8)